Four different colored flags can be hung in a row to make coded signal. How many signals can be made if a signal consists of the display of one or more flags?
Answer by ikleyn(52776)

.
4            signals if to use 1 flag only,    plus

4*3     = 12 signals if to use 2 flags only,   plus

4*3*2   = 24 signals if to use 3 flags only,   plus

4*3*2*1 = 24 signals if to use all 4 flags.


In all, 4 + 12 + 24 + 24 = 64 signals.


Answer.  64 signals.

Solved.